# DiffuBox: Refining 3D Object Detection with Point Diffusion

### v1

@@ 5 Discussion and Future Work
S2.3 Additional Qualitative Results
We present additional visualizations of box refinement over denoising steps, in both original and normalized views, in Figure S6, and visualizations of detections before and after using DiffuBox in Figure S7.
In Figure S6, we show the box refinement process of the initial noisy prediction across different denoising steps, up to 14 steps. In addition, we show the points from the normalized box view (NBV) from the perspective of the refined boxes. Observe how the points of the car (colored in black) gradually get corrected to be inside the bounding box. We visualize the performance of DiffuBox on another set of scene in Figure S7. Observe that our method is able to better align boxes after refinement, and the corrected boxes can reduce false positives, since they get corrected into the same set of points.
@@
Figure S7: Additional Qualitative Results of DiffuBox. We include additional visualizations of 3D object detection on the Lyft/Ithaca-365 dataset before and after DiffuBox’s refinement We color the ground truth boxes in green, the detector outputs in yellow, and the boxes refined with DiffuBox in blue.
S2.4 Improvement Expectations and Upper Bounds
We validate our claims that improvements in localization is the major source of performance gain in Table S11. In particular, we show that large gains in performance can be obtained —across multiple domain adaptation algorithms— if localization is correct to the ground truth position. This suggests that there is significant room for improvement in object detection performance from better localization, which DiffuBox aims to tackle.
@@
Recent domain adaptation algorithms typically rely on Statistical Normalization (SN) [47] to tackle the domain bias in object size. SN’s solution is to adapt the source dataset first, rescaling its objects (i.e. bounding box dimensions & points in bounding boxes) to match size statistics in the target domain. Afterwards, they finetune the object detector on the adapted dataset. In short, SN attempts to explicitly construct a new size prior to replace the old one. Despite great effectiveness, SN is very sensitive to the accuracy of size statistics in the target domain, and requires careful retraining for each new target domain. In contrast, our work is a more general interpretation of size normalization, and instead rely on size invariance to adapt it across all models and domains.

### v2

@@ 5 Discussion and Future Work
In addition, we report the full KITTI $\rightarrow$ Lyft performance on all three traffic participant classes with the CenterPoint [56] detector in Table S12 and DSVT [47] in Table S13. DiffuBox achieves consistent improvement across object detectors, classes, and domain adaptation methods.
Table S12: mAP for KITTI $\rightarrow$ Lyft with CenterPoint [56]. DiffuBox attains consistent improvement across object classes and domain adaptation methods.
Class Method BEV $\uparrow$ 3D $\uparrow$
0-30m 30-50m 50-80m 0-80m 0-30m 30-50m 50-80m 0-80m
Car@IoU 0.7 Direct 74.91 36.64 2.47 36.23 28.63 4.05 0.15 10.29
Direct+DiffuBox 90.25 58.65 6.95 51.38 71.02 34.91 1.48 34.40
OT 82.81 51.02 6.26 46.10 25.34 12.35 0.52 13.34
OT+DiffuBox 91.79 59.24 7.45 51.88 63.01 32.53 1.24 30.95
Pedestrian@IoU 0.5 Direct 0.85 1.22 0.02 0.49 0.41 0.47 0.01 0.19
Direct+DiffuBox 5.58 2.06 0.03 1.78 3.99 1.75 0.02 1.20
OT 1.88 2.02 0.05 0.87 0.61 1.22 0.01 0.37
OT+DiffuBox 5.78 2.32 0.03 1.78 4.04 1.90 0.02 1.16
Cyclist@IoU 0.5 Direct 23.34 0.84 0.01 8.19 16.53 0.34 0.00 5.42
Direct+DiffuBox 42.22 1.49 0.02 15.22 31.74 0.58 0.01 11.28
OT 27.40 0.95 0.01 9.73 7.32 0.05 0.01 1.98
OT+DiffuBox 42.22 1.54 0.02 15.23 31.35 0.67 0.01 11.20
Table S13: mAP for KITTI $\rightarrow$ Lyft with DSVT [47]. DiffuBox attains consistent improvement across object classes and domain adaptation methods.
Class Method BEV $\uparrow$ 3D $\uparrow$
0-30m 30-50m 50-80m 0-80m 0-30m 30-50m 50-80m 0-80m
Car@IoU 0.7 Direct 68.93 47.49 11.32 41.77 33.72 11.74 1.42 15.67
Direct+DiffuBox 89.01 63.41 17.50 55.27 65.22 36.31 5.02 35.61
OT 71.85 42.93 13.18 43.05 15.66 4.47 0.31 8.06
OT+DiffuBox 90.21 63.50 17.91 55.61 56.84 31.19 4.63 31.12
Pedestrian@IoU 0.5 Direct 16.42 6.48 0.47 8.09 11.49 4.41 0.10 5.32
Direct+DiffuBox 27.89 8.13 1.05 12.42 23.55 6.49 0.53 10.28
OT 20.89 7.47 1.06 10.65 14.83 5.46 0.33 6.84
OT+DiffuBox 27.68 7.84 1.25 12.18 25.67 6.66 0.68 10.64
Cyclist@IoU 0.5 Direct 38.46 2.04 0.00 19.82 30.16 1.56 0.00 15.65
Direct+DiffuBox 54.41 3.19 0.00 28.01 47.57 1.94 0.00 24.06
OT 43.43 2.18 0.00 22.30 14.65 1.23 0.00 8.01
OT+DiffuBox 55.26 3.04 0.00 28.75 38.12 1.89 0.00 19.52
S3.3 Additional Qualitative Results
We present additional visualizations of box refinement over denoising steps, in both original and normalized views, in Figure S8. We show the box refinement process of the initial noisy prediction across different denoising steps, up to 14 steps. In addition, we show the points from the normalized box view (NBV) from the perspective of the refined boxes. Observe how the points of the car (colored in black) gradually get corrected to be inside the bounding box.
We present additional visualizations of detections before and after using DiffuBox on another set of scene in Figure S9, and on different classes of traffic participants in Figure S10. Observe that DiffuBox’s refinement is able to better align the boxes, and the corrected boxes can reduce false positives as they get corrected into the same set of points.
@@
Figure S9: Additional Qualitative Results of DiffuBox. We include additional visualizations of 3D object detection on Lyft/Ithaca365 before and after DiffuBox’s refinement. We color the ground truth boxes in green, the detector outputs in yellow, and the boxes refined with DiffuBox in blue.
Figure S10: Additional Qualitative Results of DiffuBox for other classes. We include additional visualizations of 3D object detection for different classes. We color the ground truth boxes in green, the detector outputs in yellow, and the boxes refined with DiffuBox in blue.
S3.4 Improvement Expectations and Upper Bounds
We validate our claims that improvements in localization is the major source of performance gain in Table S14. In particular, we show that large gains in performance can be obtained – across multiple domain adaptation algorithms – if localization is corrected to the ground truth position. This suggests that there is significant room for improvement in object detection performance from better localization, which DiffuBox aims to tackle.
@@
Recent domain adaptation algorithms typically rely on Statistical Normalization (SN) [48] to tackle the domain bias in object size. SN’s solution is to adapt the source dataset first, rescaling its objects (i.e. bounding box dimensions & points in bounding boxes) to match size statistics in the target domain. Afterwards, they finetune the object detector on the adapted dataset. In short, SN attempts to explicitly construct a new size prior to replace the old one. Despite great effectiveness, SN is very sensitive to the accuracy of size statistics in the target domain, and requires careful retraining for each new target domain. In contrast, our work is a more general interpretation of size normalization, and instead rely on shape invariance to adapt it across all models and domains.